# Behavior-Regularized Diffusion Policy Optimization for Offline Reinforcement Learning

### v1

@@ Appendix C Proof for the Main Results
for each $a^{n}$ yields
$p^{\nu,s}_{n-1|n}(a^{n-1}|a^{n})=\frac{\mathbb{E}_{a^{0}\sim\nu(\cdot|s)}[q_{n|0}(a^{n}|a^{0})q_{n-1|0,n}(a^{n-1}|a^{0},a^{n})]}{\mathbb{E}_{a^{0}\sim\nu(\cdot|s)}[q_{n|0}(a^{n}|a^{0})]}.$
Eq. (24)) then follows trivially.
Then, we will show that the equality $p^{*,s}_{n-1|n}=q^{s,\pi^{*}}_{n-1|n}$ holds for
$\pi^{*}(a|s)=\frac{\nu(a|s)\exp(Q(s,a)/\eta)}{\mathbb{E}_{a\sim\nu(\cdot|s)}[\exp(Q(s,a)/\eta)]}.$
Note that Eq. (13) on every state $s$ constitutes a standard maximum entropy reinforcement learning problem (Haarnoja et al., 2018; Wu et al., 2019; Garg et al., 2023) with ending reward $Q(s,a^{0})$ and no intermediate rewards. In such cases, the optimal diffusion value function satisfies the following recursion,
$V^{*,s}_{0}(a^{0})=Q(s,a^{0})$
$V_{n}^{*,s}(a^{n})=\eta\log\mathbb{E}_{a^{n-1}\sim p^{\nu,s}_{n-1|n}(\cdot|a^{n})}\left[\exp(V_{n-1}^{*,s}(a^{n-1})/\eta)\right].$
Expanding the recursion, we have
$V_{n}^{*,s}(a^{n})=\eta\log\mathbb{E}_{a^{0}\sim p^{\nu,s}_{0|n}(\cdot|a^{n})}\left[\exp(Q(s,a^{0})/\eta)\right].$ (26)
The optimal policy of the maximum entropy reinforcement learning problem is a Boltzmann distribution defined as follows,
$p^{*,s}_{n-1|n}(a^{n-1}|a^{n})=\frac{p_{n-1|n}^{\nu,s}(a^{n-1}|a^{n})\exp(V_{n-1}^{*,s}(a^{n-1})/\eta)}{\mathbb{E}_{a^{n-1}\sim p^{\nu,s}_{n-1|n}(\cdot|a^{n})}\left[\exp(V_{n-1}^{*,s}(a^{n-1})/\eta)\right]}.$
Substituting $p_{n-1|n}^{\nu,s}$ and $V_{n-1}^{*,s}$ with Equations (24) and (26), we get
$p^{*,s}_{n-1|n}(a^{n-1}|a^{n})=\frac{q^{\nu,s}_{n-1|n}(a^{n-1}|a^{n})\mathbb{E}_{a^{0}\sim q^{\nu,s}_{0|n-1}(\cdot|a^{n-1})}\left[\exp(Q(s,a^{0})/\eta)\right]}{\mathbb{E}_{a^{0}\sim q_{0|n}^{\nu,s}(\cdot|a^{n})}\left[\exp(Q(s,a^{0}))/\eta\right]},$ (27)
where
$\mathbb{E}_{a^{0}\sim q_{0|n}^{\nu,s}(\cdot|a^{n})}\left[\exp(Q(s,a^{0}))/\eta\right]=\int q^{\nu,s}_{n-1|n}(a^{n-1}|a^{n})\mathbb{E}_{a^{0}\sim q^{\nu,s}_{0|n-1}(\cdot|a^{n-1})}\left[\exp(Q(s,a^{0})/\eta)\right]\mathrm{d}a^{n-1}.$
Unfolding the definition of $q^{\nu,s}_{n-1|n}(a^{n-1}|a^{n})$ (Eq. (25)), we have
$p^{*,s}_{n-1|n}(a^{n-1}|a^{n})$ $=\frac{\mathbb{E}_{a^{0}\sim\pi^{*}(\cdot|s)}[q^{\pi^{*},s}_{n-1|0}(a^{n-1}|a^{0})q^{\pi^{*},s}_{n|n-1}(a^{n}|a^{n-1})]}{\mathbb{E}_{a^{0}\sim\pi^{*}(\cdot|s)}[q^{\pi^{*},s}_{n|0}(a^{n}|a^{0})]}.$
Thus, we conclude that the optimal diffusion policy $p^{*,s}_{0:N}$ matches in distribution with the forward diffusion process $q^{s,\pi^{*}}_{0:N}$ originating from the initial distribution $\pi^{*}(s)$. In other word, $\pi^{*}(a|s)=\int p^{*,s}(u_{0:N})\delta(a^{0}-a)\mathrm{d}u_{0:N}$. Since $\pi^{*}$ is the optimal solution of the KL regularized objective in Eq. (4)), the proof is completed.
@@
Let $p^{\nu}$ be the reference behavior diffusion process. The optimal diffusion policy $p^{*}$ of the pathwise KL-regularized RL in Eq. (11) is also the optimal policy $\pi^{*}$ of the KL regularized RL problem in Eq. (1) in the sense that $\pi^{*}(a|s)=\int p^{*,s}_{0:N}(a^{0:N})\delta(a-a^{0})\mathrm{d}a^{0:N}\forall s\in\mathcal{S}$, where $\delta$ is the Dirac delta function.
@@
Denote the optimal state-action value function and state value functions of the pathwise KL-regularized RL as $Q^{*}$ and $V^{*}$, respectively. According to Theorem C.1, for any state $s\in\mathcal{S}$, $p^{*,s}$ is the reverse process of the forward diffusion process $q^{\pi^{*},s}$. Since the $p^{\nu,s}$ is the reverse process of $q^{\nu,s}$, we have
$\operatorname{KL}\left[p^{*,s}_{0:N}\|p^{\nu,s}_{0:N}\right]=\operatorname{KL}\left[q^{\pi^{*},s}_{0:N}\|q^{\nu,s}_{0:N}\right].$
Notice that $q^{\pi^{*},s}_{0:N}$ and $q^{\nu,s}_{0:N}$ share the same transition kernel and differ only on the initial distribution. Hence, conditioned on the initial action $a^{0}$, they match in distribution,
$\operatorname{KL}\left[q_{1:N|0}^{\pi^{*},s}(a^{0})\|q_{1:N|0}^{\nu,s}(a^{0})\right]=\mathbb{E}\left[\log\frac{\prod_{n=1}^{N}q(a^{n}|a^{n-1})}{\prod_{n=1}^{N}q(a^{n}|a^{n-1})}\right]=0,\forall a^{0}\in\mathcal{A}.$
Therefore, the KL divergence between diffusion processes $q^{\pi^{*},s}_{0:N}$ and $q_{0:N}^{\nu,s}$ equals the KL divergence between $\pi^{*}(s)$ and $\nu(s)$,
$\operatorname{KL}\left[p^{*,s}_{0:N}\|p^{\nu,s}_{0:N}\right]=\operatorname{KL}\left[q^{\pi^{*},s}_{0:N}\|q^{\nu,s}_{0:N}\right]=\operatorname{KL}\left[\pi^{*}(s)\|\nu(s)\right]+\mathbb{E}\left[\operatorname{KL}\left[q_{1:N|0}^{\pi^{*},s}\|q_{1:N|0}^{\nu,s}\right]\right]=\operatorname{KL}\left[\pi^{*}(s)\|\nu(s)\right].$
We conclude the proof by showing that the optimal soft value functions $Q^{*}$ and $V^{*}$ of the optimal diffusion policy satisfy the Bellman optimality equations of the KL regularized objective in Eq. (1),
$V^{*}(s)=\mathbb{E}_{a\sim p^{*,s}_{0}}\left[Q^{*}(s,a)\right]+\eta\operatorname{KL}\left[p^{*,s}_{0}\|\nu(s)\right]$
$Q^{*}(s,a)=R(s,a)+\gamma\mathbb{E}_{s^{\prime}\sim T(s,a)}[V^{*}(s^{\prime})].$
Here, the second equality is trivially satisfied. The first equality is satisfied because we have $\operatorname{KL}\left[p_{0}^{*,s}\|\nu(s)\right]=\operatorname{KL}\left[\pi^{*}(s)\|\nu(s)\right]=\operatorname{KL}\left[p_{0:N}^{*,s}\|p_{0:N}^{\nu,s}\right]$, and
$V^{*}(s)=\mathbb{E}_{a\sim p^{*,s}_{0}}\left[Q^{*}(s,a)\right]+\eta\operatorname{KL}\left[p_{0:N}^{*,s}\|p_{0:N}^{\nu,s}\right]$
holds by definition. ∎
@@
The proof is essentially the same as that presented in Haarnoja et al. (2018) except that we employ a bounded probability ratio to ensure the boundedness of the Q value function rather than requiring the action space to be a finite set. ∎
Proposition C.4 (Policy Improvement (adapted from Haarnoja et al. (2018))).
Let $p^{\pi_{\textrm{new}}}$ be the optimizer of the optimization problem defined in Eq. (15). Then, $V^{\pi_{\textrm{new}},s}_{n}(a)\geq V^{\pi_{\textrm{old}},s}_{n}(a)$ under Assumption 4.3.
@@
The optimizer $p^{\pi_{\textrm{new}},s,a^{n}}_{n-1|n}$ of the optimization problem Eq. (15) is guaranteed to satisfy
$-\eta\ell^{\pi_{\textrm{new}},s}_{n}(a^{n})+\underset{p^{\pi_{\textrm{new}},s,a^{n}}_{n-1|n}}{\mathbb{E}}\left[V^{\pi_{\textrm{old}},s}_{n-1}(a^{n-1})\right]\geq-\eta\ell^{\pi_{\textrm{old}},s}_{n}(a^{n})+\underset{p^{\pi_{\textrm{old}},s,a^{n}}_{n-1|n}}{\mathbb{E}}\left[V^{\pi_{\textrm{old}},s}_{n-1}(a^{n-1})\right]=V^{\pi_{\textrm{old}},s}_{n}(a^{n}).$
Recursively applying the above inequality to $V^{\pi_{\textrm{old}},s}_{n}(a^{n})=-\eta\ell^{\pi_{\textrm{old}},s}_{n}(a^{n})+\underset{p^{\pi_{\textrm{old}},s,a^{n}}_{n-1|n}}{\mathbb{E}}\left[V^{\pi_{\textrm{old}},s}_{n-1}(a^{n-1})\right]$, we get
$V^{\pi_{\textrm{old}},s}_{n}(a^{n})$ $\leq-\eta\ell^{\pi_{\textrm{old}},s}_{n}(a^{n})+\underset{p^{\pi_{\textrm{old}},s,a^{n}}_{n-1|n}}{\mathbb{E}}\left[-\eta\ell^{\pi_{\textrm{new}},s}_{n-1}(a^{n-1})+\underset{p^{\pi_{\textrm{new}},s,a^{n-1}}_{n-2|n-1}}{\mathbb{E}}\left[V^{\pi_{\textrm{old}},s}_{n-2}(a^{n-2})\right]\right]$
$\leq\underset{p^{\pi_{\textrm{old}},s,a^{n}}_{0:n-1|n}}{\mathbb{E}}\left[-\eta\sum_{i=1}^{n}\ell^{\pi_{\textrm{new}},s}_{i}(a^{i})+Q^{\pi_{\textrm{old}}}(s,a^{0})\right]$
$\leq\underset{p^{\pi_{\textrm{old}},s,a^{n}}_{0:n-1|n}}{\mathbb{E}}\left[-\eta\sum_{i=1}^{n}\ell^{\pi_{\textrm{new}},s}_{i}(a^{i})+R(s,a^{0})+\gamma\mathbb{E}_{s^{\prime}\sim T(s,a^{0})}\left[-\eta\sum_{i=1}^{N}\ell_{i}^{\pi_{\textrm{new}},s^{\prime}}(a^{\prime i})+Q^{\pi_{\textrm{old}}}(s,a^{\prime 0})\right]\right]$
[…]
$\leq V_{n}^{\pi_{\textrm{new}},s}(a^{n}),$
where the convergence to $V_{n}^{\pi_{\textrm{new}},s}$ is guaranteed by the Lemma C.3. ∎

### v2

@@ Appendix C Proof for the Main Results
The corresponding Lagrange function is
$L(p^{\nu,s}_{n-1|n},\lambda)$ $=\mathbb{E}_{a^{0}\sim\nu(\cdot|s)}\left[q_{n|0}(a^{n}|a^{0})D_{\rm KL}\left[q_{n-1|0,n}(\cdot|a^{0},a^{n})\|p_{n-1|n}^{\nu,s}(\cdot|a^{n})\right]\right]+\lambda(1-\int p_{n-1|n}^{\nu,s}(a^{n-1}|a^{n})\mathrm{d}a^{n-1}),$ (25)
where $\lambda$ is the multiplier. Setting $\frac{\partial L}{\partial p_{n|n-1}^{\nu,s}}=0$, we obtain
$p^{\nu,s}_{n-1|n}(a^{n-1}|a^{n})$ $=-\frac{1}{\lambda}\mathbb{E}_{a^{0}\sim\nu^{0}(\cdot|s)}\left[q_{n|0}(a^{n}|a^{0})q_{n-1|0,n}(a^{n-1}|a^{0},a^{n})\right].$ (26)
Since $\int p_{n-1|n}^{\nu,s}(a^{n-1}|a^{n})\mathrm{d}a^{n-1}=1$, it follows that
$\lambda$ $=-\int\mathbb{E}_{a^{0}\sim\nu(\cdot|s)}\left[q_{n|0}(a^{n}|a^{0})q_{n-1|0,n}(a^{n-1}|a^{0},a^{n})\right]\mathrm{d}a^{n-1}$ (27)
$=-\mathbb{E}_{a^{0}\sim\nu(\cdot|s)}\left[q_{n|0}(a^{n}|a^{0})\right].$
Substituting Eq. (27) into Eq. (26),
$p^{\nu,s}_{n-1|n}(a^{n-1}|a^{n})$ $=\frac{\mathbb{E}_{a^{0}\sim\nu(\cdot|s)}\left[q_{n|0}(a^{n}|a^{0})q_{n-1|0,n}(a^{n-1}|a^{0},a^{n})\right]}{\mathbb{E}_{a^{0}\sim\nu^{0}(\cdot|s)}\left[q_{n|0}(a^{n}|a^{0})\right]}$ (28)
$=\frac{\mathbb{E}_{a^{0}\sim\nu(\cdot|s)}\left[q_{n|0}(a^{n}|a^{0})q_{n-1|0,n}(a^{n-1}|a^{0},a^{n})\right]}{q_{n}(a^{n})}$
$=\int\frac{v(a^{0})q_{n|0}(a^{n}|a^{0})q_{n-1|0,n}(a^{n-1}|a^{0},a^{n})}{q_{n}(a^{n})}\mathrm{d}a^{0}$
$=q_{n-1|n}^{\nu,s}(a^{n-1}|a^{n}),$
That is, through optimizing Eq. (24), the derived behavior diffusion policy is consistent with $q_{n-1|n}^{\nu,s}$ from the forward process.
Next, we prove that the optimizer of Eq. (14) also satisfies $p^{*,s}_{n-1|n}=q^{\pi_{Q},s}_{n-1|n}$.
To do this, we first recognize that the reverse (generative) process of the diffusion policy constitutes a Markov Decision Process (MDP), where the state at time step $n$ is the tuple $(s,a^{n})$ and the agent’s policy follows $a^{n-1}\sim p^{\pi,s}_{n-1|n}(\cdot|a^{n})$. The transition in this MDP is implicit and deterministic: upon selecting the action $a^{n-1}$, the state immediately transitions from $(s,a^{n})$ to $(s,a^{n-1})$. At the last time step $n=0$, the agent will receive an ending reward of $Q(s,a^{0})$. In this sense, the problem defined in Eq. (14) can be recognized as a behavior-regularized RL problem with Gaussian parameterized policy $p^{\pi,s}_{n-1|n}$ and behavior policy $p^{\nu,s}_{n-1|n}$. Following Wu et al. (2019), the optimal diffusion value functions satisfy the recursion:
$V_{0}^{*,s}(a^{0})$ $=Q(s,a^{0}),$ (29)
$V_{n}^{*,s}(a^{n})$ $=\eta\log\mathbb{E}_{a^{n-1}\sim p^{\nu,s}_{n-1|n}}\left[\exp(V^{*,s}_{n-1}(a^{n-1})/\eta)\right],$
and the optimal diffusion policy is given by:
$p^{*,s}_{n-1|n}(a^{n-1}|a^{n})=\frac{p^{\nu,s}_{n-1|n}(a^{n-1}|a^{n})\exp(V^{*,s}_{n-1}(a^{n-1})/\eta)}{\mathbb{E}_{a^{n-1}\sim p^{\nu,s}_{n-1|n}}\left[\exp(V^{*,s}_{n-1}(a^{n-1})/\eta)\right]}.$ (30)
Expanding the recursion of diffusion value functions, we have
$V^{*,s}_{n}(a^{n})$ $=\eta\log\mathbb{E}_{a^{n-1}\sim p^{\nu,s}_{n-1|n}}\left[\exp(V^{*,s}_{n-1}(a^{n-1})/\eta)\right]$ (31)
$=\eta\log\mathbb{E}_{a^{n-1}\sim p^{\nu,s}_{n-1|n},a^{n-2}\sim p^{\nu,s}_{n-2|n-1}}\left[\exp(V^{*,s}_{n-2}(a^{n-2})/\eta)\right]$
=$\ldots$= …
$=\eta\log\mathbb{E}_{a^{0}\sim p^{\nu,s}_{0|n}}\left[\exp(V^{*,s}_{0}(a^{0})/\eta)\right]$
$=\eta\log\mathbb{E}_{a^{0}\sim p^{\nu,s}_{0|n}}\left[\exp(Q(s,a^{0})/\eta)\right].$
Similarly, the optimal policy can be expressed as:
$p^{*,s}_{n-1|n}(a^{n-1}|a^{n})$ $=\frac{p^{\nu,s}_{n-1|n}(a^{n-1}|a^{n})\mathbb{E}_{a^{0}\sim p^{\nu,s}_{0|n-1}}\left[\exp(Q(s,a^{0})/\eta)\right]}{\mathbb{E}_{a^{n-1}\sim p^{\nu,s}_{n-1|n}}\left[\mathbb{E}_{a^{0}\sim p^{\nu,s}_{0|n-1}}\left[\exp(Q(s,a^{0})/\eta)\right]\right]}.$ (32)
Let’s consider the following expression:
$\mathbb{E}_{a^{0}\sim p^{\nu,s}_{0|n}}\left[\exp(Q(s,a^{0})/\eta)\right]$ $=\int p^{\nu,s}_{0|n}(a^{0}|a^{n})\exp(Q(s,a^{0})/\eta)\mathrm{d}a^{0}$ (33)
$=\int q^{\nu,s}_{0|n}(a^{0}|a^{n})\exp(Q(s,a^{0})/\eta)\mathrm{d}a^{0}$
$=\int\frac{q^{\nu,s}_{n|0}(a^{n}|a^{0})\nu(a^{0}|s)}{q^{\nu,s}_{n}(a^{n})}\exp(Q(s,a^{0})/\eta)\mathrm{d}a^{0}$
$=\frac{1}{Z(s)}\int\frac{q^{\nu,s}_{n|0}(a^{n}|a^{0})\pi_{Q}(a^{0}|s)}{q^{\nu,s}_{n}(a^{n})}\mathrm{d}a^{0}$
$=\frac{1}{Z(s)q^{\nu,s}_{n}(a^{n})}\mathbb{E}_{a^{0}\sim\pi_{Q}}\left[q^{\nu,s}_{n|0}(a^{n}|a^{0})\right],$
where $Z(s)$ is the normalizing factor, and the second equation follows from $p^{\nu,s}_{n-1|n}=q^{\nu,s}_{n-1|n}$. Substituting $\mathbb{E}_{a^{0}\sim p^{\nu,s}_{0|n}}\left[\exp(Q(s,a^{0})/\eta)\right]$ and $\mathbb{E}_{a^{0}\sim p^{\nu,s}_{0|n-1}}\left[\exp(Q(s,a^{0})/\eta)\right]$ with Eq. (33) into Eq. (32),
$p^{*,s}_{n-1|n}(a^{n-1}|a^{n})$ $=\frac{p^{\nu,s}_{n-1|n}(a^{n-1}|a^{n})q^{\nu,s}_{n}(a^{n})\mathbb{E}_{a^{0}\sim\pi_{Q}}\left[q^{\nu,s}_{n-1|0}(a^{n-1}|a^{0})\right]}{q^{\nu,s}_{n-1}(a^{n-1})\mathbb{E}_{a^{0}\sim\pi_{Q}}\left[q^{\nu,s}_{n|0}(a^{n}|a^{0})\right]}$ (34)
$=\frac{q^{\nu,s}_{n-1|n}(a^{n-1}|a^{n})q^{\nu,s}_{n}(a^{n})\mathbb{E}_{a^{0}\sim\pi_{Q}}\left[q^{\nu,s}_{n-1|0}(a^{n-1}|a^{0})\right]}{q^{\nu,s}_{n-1}(a^{n-1})\mathbb{E}_{a^{0}\sim\pi_{Q}}\left[q^{\nu,s}_{n|0}(a^{n}|a^{0})\right]}$
$=\frac{q^{\nu,s}_{n|n-1}(a^{n}|a^{n-1})\mathbb{E}_{a^{0}\sim\pi_{Q}}\left[q^{\nu,s}_{n-1|0}(a^{n-1}|a^{0})\right]}{\mathbb{E}_{a^{0}\sim\pi_{Q}}\left[q^{\nu,s}_{n|0}(a^{n}|a^{0})\right]}$
$=\frac{q^{\pi_{Q},s}_{n|n-1}(a^{n}|a^{n-1})\mathbb{E}_{a^{0}\sim\pi_{Q}}\left[q^{\pi_{Q},s}_{n-1|0}(a^{n-1}|a^{0})\right]}{\mathbb{E}_{a^{0}\sim\pi_{Q}}\left[q^{\pi_{Q},s}_{n|0}(a^{n}|a^{0})\right]}$
$=q^{\pi_{Q},s}_{n-1|n}(a^{n-1}|a^{n}),$
where the second line follows from $p^{\nu,s}_{n-1|n}=q^{\nu,s}_{n-1|n}$ and the fourth line follows from the fact that $q^{\nu,s}_{i|j}=q^{\pi_{Q},s}_{i|j}$ for any $i>j$.
In conclusion, we obtain $p^{*,s}_{n-1|n}(a^{n-1}|a^{n})=q^{\pi_{Q},s}_{n-1|n}(a^{n-1}|a^{n})$, meaning that the reverse process $p^{*,s}$ exactly corresponds to the forward process starting from $\pi_{Q}(a|s)\propto\nu(a|s)\exp(Q(s,a)/\eta)$.
@@
Let $p^{\nu}$ be the behavior diffusion process. The optimal diffusion policy $p^{*}$ of the pathwise KL-regularized RL problem in Eq. (12) is also the optimal policy $\pi^{*}$ of the KL regularized objective in Eq. (1), in the sense that $\pi^{*}(a|s)=\int p^{*,s}_{0:N}(a^{0:N})\delta(a-a^{0})\mathrm{d}a^{0:N}\ \forall s\in\mathcal{S}$, where $\delta$ is the Dirac delta function.
@@
Denote the optimal Q-value function of the pathwise KL-regularized RL as $Q^{*}$. According to Theorem C.1, for any state $s\in\mathcal{S}$, $p^{*,s}$ is the reverse process of the forward diffusion process $q^{\pi_{Q^{*}},s}$. Since the $p^{\nu,s}$ is also the reverse process of $q^{\nu,s}$, we have
$D_{\rm KL}\left[p^{*,s}_{0:N}\|p^{\nu,s}_{0:N}\right]=D_{\rm KL}\left[q^{\pi_{Q^{*}},s}_{0:N}\|q^{\nu,s}_{0:N}\right].$ (35)
Notice that $q^{\pi_{Q^{*}},s}_{0:N}$ and $q^{\nu,s}_{0:N}$ share the same transition kernel and differ only in the initial distribution. Hence,
$D_{\rm KL}\left[p^{*,s}_{0:N}\|p^{\nu,s}_{0:N}\right]$ $=D_{\rm KL}\left[q^{\pi_{Q^{*}},s}_{0:N}\|q^{\nu,s}_{0:N}\right]$ (36)
$=D_{\rm KL}\left[q^{\pi_{Q^{*}},s}_{0}\|q^{\nu,s}_{0}\right]+D_{\rm KL}\left[q^{\pi_{Q^{*}},s}_{1:N|0}\|q^{\nu,s}_{1:N|0}\right]$
$=D_{\rm KL}\left[q^{\pi_{Q^{*}},s}_{0}\|q^{\nu,s}_{0}\right]$
$=D_{\rm KL}\left[\pi_{Q^{*}}(\cdot|s)\|\nu(\cdot|s)\right].$
Therefore, the optimal Q-value function $Q^{*}(s,a)$ satisfies
$Q^{*}(s,a)$ $=R(s,a)+\gamma\mathbb{E}_{a^{\prime 0:N}\sim p^{\*,s^{\prime}}_{0:N}}\left[Q^{*}(s^{\prime},a^{\prime}0)-\eta D_{\rm KL}\left[p^{*,s^{\prime}}_{0:N}\|p^{\nu,s^{\prime}}_{0:N}\right]\right]$ (37)
$=R(s,a)+\gamma\mathbb{E}_{a^{\prime}\sim\pi_{Q^{*}}}\left[Q^{*}(s^{\prime},a^{\prime})-\eta D_{\rm KL}\left[\pi_{Q^{*}}(\cdot|s^{\prime})\|\nu(\cdot|s^{\prime})\right]\right]$
$=R(s,a)+\gamma\eta\log\mathbb{E}_{a^{\prime}\sim\nu(\cdot|s^{\prime})}\left[\exp(Q^{*}(s^{\prime},a^{\prime})/\eta)\right],$
which is exactly the optimal Bellman iteration of the problem defined in Eq. (1). This means that $Q^{*}$ is also the optimal Q-value function of the problem defined in Eq. (1), and $\pi_{Q^{*}}$ is correspondingly the optimal solution $\pi^{*}$. Using Theorem C.1, we know that $p^{*,s}$ and $\pi^{*}$ are equivalent:
$\pi^{*}(a|s)=\int p^{*,s}_{0:N}(a^{0:N})\delta(a-a^{0})\mathrm{d}a^{0:N}\quad\forall s\in\mathcal{S}.$ (38)
The proof is completed. ∎
@@
Define the KL-augmented reward as
$\tilde{R}^{\pi}(s,a)\coloneqq R(s,a)-\eta\mathbb{E}_{a^{\prime 0:N}\sim p^{\pi,s^{\prime}}_{0:N}}\left[\sum_{n=1}^{N}D_{\rm KL}\left[p_{n-1|n}^{\pi,s^{\prime}}\|p_{n-1|n}^{\nu,s^{\prime}}\right]\right],$
and rewrite the policy evaluation update rule as
$Q^{k+1}(s,a)\leftarrow\tilde{R}^{\pi}(s,a)+\gamma\mathbb{E}_{a^{\prime 0:N}\sim p^{\pi,s^{\prime}}_{0:N}}\left[Q^{k}(s^{\prime},a^{\prime 0})\right].$
When $\gamma\in(0,1)$, this update satisfies the $\gamma$-contraction property, which is demonstrated in Haarnoja et al. (2018) to converge to a unique solution. Since $Q^{\pi}$ satisfies $Q^{\pi}=\mathcal{B}^{\pi}Q^{\pi}$, we obtain that the sequence $Q^{k}$ converges to $Q^{\pi}$. Note that we employ a bounded probability ratio to ensure the boundedness of the $Q$-value function rather than requiring the action space to be a finite set. ∎
Proposition C.4 (Soft Policy Improvement).
Let $p^{\pi_{\textrm{new}}}$ be the optimizer of the problem defined in Eq. (16). Under Assumption 4.3, $V^{\pi_{\textrm{new}},s}_{n}(a^{n})\geq V^{\pi_{\textrm{old}},s}_{n}(a^{n})$ holds for all $n\in\{0,1,\ldots,N\}$ and $(s,a)\in\mathcal{S}\times\mathcal{A}$.
@@
Since $p^{\pi_{\textrm{new}},s,a^{n}}$ is the optimizer of the problem 16, it satisfies
$-\eta\ell^{\pi_{\textrm{old}},s}_{n}(a^{n})+\underset{p^{\pi_{\textrm{old}},s,a^{n}}_{n-1|n}}{\mathbb{E}}\left[V^{\pi_{\textrm{old}},s}_{n-1}(a^{n-1})\right]=V^{\pi_{\textrm{ old}},s}_{n}(a^{n})\leq-\eta\ell^{\pi_{\textrm{new}},s}_{n}(a^{n})+\underset{p^{\pi_{\textrm{new}},s,a^{n}}_{n-1|n}}{\mathbb{E}}\left[V^{\pi_{\textrm{old}},s}_{n-1}(a^{n-1})\right].$
For $n=0$, recursively applying the above inequality leads to
$V_{0}^{\pi_{\textrm{old}},s}(a^{0})$
$=Q^{\pi_{\textrm{old}}}(s,a^{0})$
$=R(s,a^{0})+\gamma\mathbb{E}_{a^{\prime 0:N}\sim p^{\pi_{\textrm{old}},s^{\prime}}_{0:N}}\left[Q^{\pi_{\textrm{old}}}(s,a^{\prime 0})-\eta\sum_{i=1}^{N}\ell_{i}^{\pi_{\textrm{old}},s^{\prime}}(a^{\prime i})\right]$
$=R(s,a^{0})+\gamma\mathbb{E}_{a^{\prime 0:N}\sim p^{\pi_{\textrm{old}},s^{\prime}}_{0:N}}\left[V^{\pi_{\textrm{old}},s^{\prime}}_{0}(a^{\prime 0})-\eta\sum_{i=1}^{N}\ell_{i}^{\pi_{\textrm{old}},s^{\prime}}(a^{\prime i})\right]$
$=R(s,a^{0})+\gamma\mathbb{E}_{a^{\prime N}\sim p^{\pi_{\textrm{old}},s^{\prime}}_{N}}\left[V^{\pi_{\textrm{old}},s^{\prime}}_{N}(a^{\prime N})\right]$
$=R(s,a^{0})+\gamma\mathbb{E}_{a^{\prime N}\sim p^{\pi_{\textrm{new}},s^{\prime}}_{N},a^{\prime N-1}\sim p^{\pi_{\textrm{old}},s^{\prime}}_{N-1|N}}\left[V^{\pi_{\textrm{old}},s^{\prime}}_{N-1}(a^{\prime N-1})-\eta\ell_{N}^{\pi_{\textrm{old}},s^{\prime}}(a^{\prime N})\right]$
$\leq R(s,a^{0})+\gamma\mathbb{E}_{a^{\prime N}\sim p^{\pi_{\textrm{new}},s^{\prime}}_{N},a^{\prime N-1}\sim p^{\pi_{\textrm{new}},s^{\prime}}_{N-1|N}}\left[V^{\pi_{\textrm{old}},s^{\prime}}_{N-1}(a^{\prime N-1})-\eta\ell_{N}^{\pi_{\textrm{new}},s^{\prime}}(a^{\prime N})\right]$
$=R(s,a^{0})+\gamma\mathbb{E}_{a^{\prime N-1:N}\sim p^{\pi_{\textrm{new}},s^{\prime}}_{N-1:N},a^{\prime N-2}\sim p^{\pi_{\textrm{old}},s^{\prime}}_{N-2|N-1}}\left[V^{\pi_{\textrm{old}},s^{\prime}}_{N-2}(a^{\prime N-2})-\eta\ell_{N-1}^{\pi_{\textrm{old}},s^{\prime}}(a^{\prime N-1})-\eta\ell_{N}^{\pi_{\textrm{new}},s^{\prime}}(a^{\prime N})\right]$
$\leq R(s,a^{0})+\gamma\mathbb{E}_{a^{\prime N-1:N}\sim p^{\pi_{\textrm{new}},s^{\prime}}_{N-1:N},a^{\prime N-2}\sim p^{\pi_{\textrm{new}},s^{\prime}}_{N-2|N-1}}\left[V^{\pi_{\textrm{old}},s^{\prime}}_{N-2}(a^{\prime N-2})-\eta\ell_{N-1}^{\pi_{\textrm{new}},s^{\prime}}(a^{\prime N-1})-\eta\ell_{N}^{\pi_{\textrm{new}},s^{\prime}}(a^{\prime N})\right]$
$\ldots$
$\leq R(s,a^{0})+\gamma\mathbb{E}_{a^{\prime 0:N}\sim p^{\pi_{\textrm{new}},s^{\prime}}_{0:N}}\left[V^{\pi_{\textrm{old}},s^{\prime}}_{0}(a^{\prime 0})-\eta\sum_{i=0}^{N}\ell_{i}^{\pi_{\textrm{new}},s^{\prime}}(a^{\prime i})\right]$
$\ldots$
$\leq V_{0}^{\pi_{\textrm{new}},s}(a^{0}),$
where the fifth line follows from the definition of the diffusion value functions, the seventh line follows from the above inequality, and the last line follows from recursively expanding the values of successor states. For $n\in\{1,2,\ldots,N\}$, the proof is similar:
$V^{\pi_{\textrm{old}},s}_{n}(a^{n})$
$=-\eta\ell^{\pi_{\textrm{old}},s}_{n}(a^{n})+\mathbb{E}_{p^{\pi_{\textrm{old}},s,a^{n}}_{n-1|n}}\left[V^{\pi_{\textrm{old}},s}_{n-1}(a^{n-1})\right]$
$\leq-\eta\ell^{\pi_{\textrm{new}},s}_{n}(a^{n})+\mathbb{E}_{p^{\pi_{\textrm{new}},s,a^{n}}_{n-1|n}}\left[V^{\pi_{\textrm{old}},s}_{n-1}(a^{n-1})\right]$
$\leq-\eta\ell^{\pi_{\textrm{new}},s}_{n}(a^{n})+\mathbb{E}_{p^{\pi_{\textrm{new}},s,a^{n}}_{n-1|n}}\left[-\eta\ell^{\pi_{\textrm{new}},s}_{n-1}(a^{n-1})+\mathbb{E}_{p^{\pi_{\textrm{new}},s,a^{n-1}}_{n-2|n-1}}\left[V^{\pi_{\textrm{old}},s}_{n-2}(a^{n-2})\right]\right]$
$\ldots$
$\leq\mathbb{E}_{p^{\pi_{\textrm{new}},s,a^{n}}_{0:n-1|n}}\left[-\eta\sum_{i=1}^{n}\ell^{\pi_{\textrm{new}},s}_{i}(a^{i})+V^{\pi_{\textrm{old}},s}_{0}(a^{0})\right]$
$\leq\mathbb{E}_{p^{\pi_{\textrm{new}},s,a^{n}}_{0:n-1|n}}\left[-\eta\sum_{i=1}^{n}\ell^{\pi_{\textrm{new}},s}_{i}(a^{i})+V^{\pi_{\textrm{new}},s}_{0}(a^{0})\right]$
$=V_{n}^{\pi_{\textrm{new}},s}(a^{n})$
where the last inequality is obtained previously. This completes the proof. Note that the $n=0$ case already guarantees the policy improvement in the environment MDP. ∎